OpenStudy (anonymous):

JACK borrowed his father's luxury car and promised to return it with a full tank of premium gas , which cost $2.50 per gallon.From experience ,he knows that he needs 15.5 gallons .But he has only $38(and no credit card),which isnt enough.He decides to get as much premium possible to fill the remainder of the tank with regular gas,which costs $2.30 per gallon.How much of each type of gas should he get? @timo86m

OpenStudy (anonymous):

2.5*x+2.3y=15.5 I think is the right equation.

OpenStudy (anonymous):

1 of them anyways

OpenStudy (anonymous):

could you go on further?

OpenStudy (anonymous):

2.5*x+2.3y=15.5 x+y=38

OpenStudy (anonymous):

2.5*x+2.3y=15.5 for how much gas he needs x+y=38 for how much cash he has.

OpenStudy (anonymous):

can u please work it out for me to see.aint so sure about my answer

OpenStudy (anonymous):

2.5*x+2.3y=15.5 x+y=38 just solve by elimination or substitution.

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

it's easy :)

OpenStudy (anonymous):

yeah it is.:-)

OpenStudy (anonymous):

think i got it wrong tho

OpenStudy (anonymous):

why

OpenStudy (anonymous):

2.5*x+2.3y= 38 x+y=15.5

OpenStudy (anonymous):

Now it is right :D

OpenStudy (anonymous):

oh okay.gonna work it out and see

OpenStudy (anonymous):

oh ok.i was right then.thanks